### v1

# Positional Prompt Tuning for Efficient 3D Representation Learning

Abstract
Point cloud analysis has achieved significant development and is well-performed in multiple downstream tasks like point cloud classification and segmentation, etc. Being conscious of the simplicity of the position encoding structure in Transformer-based architectures, we attach importance to the position encoding as a high-dimensional part and the patch encoder to offer multi-scale information. Together with the sequential Transformer, the whole module with position encoding comprehensively constructs a multi-scale feature abstraction module that considers both the local parts from the patch and the global parts from center points as position encoding. With only a few parameters, the position embedding module fits the setting of PEFT (Parameter-Efficient Fine-Tuning) tasks pretty well. Thus we unfreeze these parameters as a fine-tuning part. At the same time, we review the existing prompt and adapter tuning methods, proposing a fresh way of prompts and synthesizing them with adapters as dynamic adjustments. Our Proposed method of PEFT tasks, namely PPT, with only 1.05 of parameters for training, gets state-of-the-art results in several mainstream datasets, such as 95.01% accuracy in the ScanObjectNN OBJ_BG dataset. Codes will be released at https://github.com/zsc000722/PPT.
Introduction
With the increasing popularity of scanning devices such as RGBD cameras and liDAR, we are witnessing the rise of a new modality of data: 3D point clouds. It has a wide range of applications for many tasks, such as autonomous driving and robot grasping, etc. Because of the huge potential for these downstream tasks, technological updates in 3D point cloud representation learning are iterating very fast, from PointNet (Qi et al. 2017a), PointNet++ (Qi et al. 2017b) and then all the way down to the Transformer based methods (Pang et al. 2022; Yu et al. 2022; Zhang et al. 2022; Zhao et al. 2021; Qi et al. 2023a) which now occupying the dominant position in this field. Within all these Transformer-based 3D representation learning methods, Position Encoding plays an important role in offering location specificity for Transformer architecture which can not distinguish tokens from different positions. Different from other domains like vision or language, the position encoding in the point cloud domain usually is not designed elaborately, like the rotary position embedding(Su et al. 2024) which is now widely adopted in the NLP domain to implement relative position embedding in an absolute manner. Instead, lightweight MLP with cluster centers as input serves as the position encoding module. Noticing this circumstance, we then wonder why a simple MLP as position encoding can efficiently facilitate the performance in 3D representation learning?
We hold the opinion that the position encoding MLP taking center points as input, together with the patch encoder which deals with the neighbor points around the cluster centers, compose a multi-scale information extractor. Then the multi-scale patch embeddings of points are fed into the Transformer for feature subsequent feature extraction. The Max or Mean Pooling operation doesn’t break the disordered and continuous property. Thus this multi-scale method which focuses on both local and global features gains excellent performance in 3D representation learning. As shown in Figure 1(b), the average attention distance of position encoding patterns is relatively centralized while the local patch patterns hold a dispersed attention distribution, indicating the diverse concern levels of position encoding and patch tokens. In the meantime, the position encoding itself only needs a few parameters (0.2), so it is suitable for promoting the performance of PEFT (Parameter-Efficient Fine-tuning) tasks.
PEFT methods usually freeze most parameters of pre-trained models, only a few selected parameters and some other ones inserted are trainable during tuning, therefore significantly releasing the demand for computational resources of full fin-tuning. In the meantime by only partly saving the trainable parameters, the pressure of storage space is also released. For language or vision models, there are Adapter tuning (Houlsby et al. 2019; Li et al. 2023), Prompt tuning (Lester, Al-Rfou, and Constant 2021; Jia et al. 2022a), and LoRA (Hu et al. 2021), which introduces extra trainable layers or prompts into pre-trained models and gets results comparable or even better than full fine-tuning. So it is natural to implement the same approach on 3D point cloud models. We attach importance to the positional embedding of point cloud representation as well as fine-tuning, unfreezing the position encoding part in the fine-tuning process, and thus get better results.
In the meantime, we draw our attention to plenty of different prompt-based PEFT methods in the point cloud analysis domain (Zhou et al. 2024; Zha et al. 2023), which shares a similar paradigm. Each one of these methods inserts prompts into the encoder inputs. Instead of static prompt tokens that are selected manually, these prompts are dynamically generated by elaborate structures like the DGCNN layer in IDPT, or linear layers followed by activate function in DAPT. Both of them use an adapter-like way to generate extra prompts from the Transformer layer outputs, which brings them satisfactory results. Nevertheless, considering the prompts in the NLP domain, they are linguistically meaningful, like a series of adjectives depicting the fine-grained characteristic of the object or a specific description for the exact question. The aforementioned adapter-generated prompts are more likely to aggregate the information, thus there is nothing new for the model to get from the prompts. Nevertheless, our PPT adopts encoded sampling centers, together with the trainable positional embedding as extra prompt tokens, as shown in Figure 2. There is a fundamental distinction between this form of prompts and the previously discussed aggregate-based prompts. These prompts, obtained through sampling and clustering methods, inherently carry physical meaning in real space: specifically, the three-dimensional positional information of points. By inputting such prompts into subsequent networks and dynamically adjusting them through adapters between the Transformer layers, we can better generalize the information from pre-trained encoders to downstream tasks during training.
Our entire network architecture is remarkably straightforward and can be seamlessly integrated into transferring tasks of any Transformer-based point cloud pre-trained model. This indicates our structure is highly reusable and holds significant potential for improvement.
Considering these perspectives, we propose our Positional Prompt Tuning, namely PPT. With only a few trainable parameters, we emphasize the importance of positional embedding layers and set them trainable, so as the initial encoder which transfers point clouds into patch tokens. We also insert simple trainable adapter layers between each layer of the Transformer encoder to adjust the weighted features dynamically.
Our main contributions can be summarized as follows:
We emphasize the significance of positional embedding in 3D point cloud representation learning, and conduct extensive experiments to validate its effectiveness.
We revisit the Parameter-efficient Prompt Tuning problem in a Prompt and Adapter manner, and propose a quite simple method, Positional Prompt Tuning (PPT), combining the above two points.
With only 5 of the trainable parameters compared with the full fine-tuning, our PPT significantly reduces the demand for storage space during fine-tuning. Meanwhile, extensive experiments have also shown that our PPT substantially reduces the time of fine-tuning while achieving on par or even higher performance, e.g., 1.26 accuracy increasing on ReCon under ModelNet 1k point, and 4.82 on Point-MAE under ScanObjectNN dataset.
Related Works
3D Representation Learning
3D Representation Learning includes point-based (Qi et al. 2017a, b), voxel-based (Maturana and Scherer 2015), and multiview-based methods (Su et al. 2015; Hamdi, Giancola, and Ghanem 2021), etc. Due to the sparse but geometry-informative representation, point-based methods (Qian et al. 2022; Engel, Belagiannis, and Dietmayer 2021) have become mainstream approaches in object classification (Wu et al. 2015; Uy et al. 2019). Voxel-based CNN methods (Yi et al. 2017; Deng et al. 2021) provide dense representation and translation invariance, achieving outstanding performance in object detection (Dai et al. 2017) and segmentation (Yi et al. 2016; Armeni et al. 2016). Furthermore, due to the vigorous development of attention mechanisms (Vaswani et al. 2017), 3D Transformers (Engel, Belagiannis, and Dietmayer 2021; Liu et al. 2021; Mao et al. 2021) have also brought about effective representations for downstream tasks. Recently, 3D self-supervised representation learning has been widely studied. PointContrast (Xie et al. 2020) leverages contrastive learning across different views to acquire discriminative 3D scene representations. Point-BERT (Yu et al. 2022) and Point-MAE (Pang et al. 2022) first introduce masked modeling pretraining into 3D. ACT (Dong et al. 2023) pioneers cross-modal geometry understanding via 2D/language foundation models. ReCon (Qi et al. 2023a, 2024) further proposes to unify generative and contrastive learning. Facilitated by foundation vision-language models like CLIP (Radford et al. 2021), another line of works are proposed towards open-world 3D representation learning (Xue et al. 2023; Peng et al. 2023; Zhang, Dong, and Ma 2023; Zhang et al. 2023a; Ding et al. 2023; Fan et al. 2023; Qi et al. 2023b).
Parameter-Efficient Fine-Tuning
Pre-training and fine-tuning paradigm has become a mainstream approach in large models to maximize their semantic understanding ability for solving downstream tasks in multiple subdivisions. However, the full fine-tuning paradigm used in most cases consumes considerable storage, and may also lead to catastrophic forgetting, degrading model performance. To solve this problem, researchers in the field of NLP and 2D have proposed many effective methods as PEFT (Parameter-Efficient Fine-tuning). Prompt Tuning (Lester, Al-Rfou, and Constant 2021) and Prefix tuning (Li and Liang 2021) introduce extra inputs or vectors as the only trainable parts, while Adapter tuning (Houlsby et al. 2019) inserts additional modules between layers as the trainable part. LoRA (Hu et al. 2021) adopts low-rank approximation matrixes in a parallel manner to simulate the full fine-tuning process. VPT (Jia et al. 2022b) then introduces prompt tuning modules into image pre-trained models. These methods have achieved substantial achievements in their own specialty. So far in the 3D domain, several methods discussing the PEFT on point cloud pre-train models demonstrate their effectiveness. IDPT (Zha et al. 2023) generates instance-aware dynamic prompts instead of static prompts with a DGCNN module. DAPT (Zhou et al. 2024) then caps every layer with a light-weight TFTS layer and uses dynamic adapters to generate prompts, internally inserts in the inputs of encoders. Point-PEFT (Tang et al. 2024) constructs a point-prior bank to store feature templates of training data and uses a parameter-free attention mechanism for feature aggregation to generate prompts into first encoder layers.
PPT: Positional Prompt Tuning
The existing methods of PEFT mostly summarize and transfer some existing ways from other fields to the current domain, and improve them to get better results. We attach importance to positional embedding in 3D missions and emphasize the role of position information in 3D tasks and then revisit the PEFT problem from the perspective of prompts and adapters. With several simple modifications of the network, we propose our PPT, Positional Prompt Tuning method which comprehensively considers both two aspects above, and achieves excellent results. In this section, we discuss the specifics of our PPT as follows.
Why does positional prompt matter?
Positional encoding has always been a vital part of all Transformer-based structures, for their input form naturally lacks positional information. Unlike the comprehensive research in NLP (Su et al. 2024) and 2D vision (Wu et al. 2021) field, the positional encoding pattern is not as well-designed for the point cloud domain. Mostly in the 3D domain, the positional information is aggregated by passing patch centers into a simple MLP and added to the input patch tokens (Pang et al. 2022; Yu et al. 2022). Since the input of the position encoding layer is the cluster centers of the target point cloud, this highly semantic-rich input form can achieve excellent results barely with a simple MLP layer, even better than some commonly used position encoding forms in NLP, such as sinusoidal position encoding or RoPE. In fact, after re-examining the location information in 3D coped with MLP, we find that this form of position encoding is actually a multi-scale feature combined with the information of patch tokens, or to say, is a little PointNet. Due to the disorder property of point cloud, the current mainstream point cloud feature extraction network, such as PointNet (Qi et al. 2017a), generally adopts the following structure when a point cloud set is given:
Where usually represents MLP or convolutional layer as point level feature extractor and denotes max or mean pooling layer. Due to the disordered and continuous characteristics of point clouds, a sequence-independent feature aggregation layer like max or mean pooling is usually adopted as adding and maximizing don’t break this property. In the point Transformer-based methods, a group of centers’ coordinates are embedded and then added to patch token inputs as positional information. With center points selected by methods like FPS and their neighbor points are given, their basic structure are as follows:
where denotes the Transformer Encoder as the feature extractor, and denotes MLP encoder layer. We separately investigate the two parts of the input of the encoder which are originally added together.
The branch accepts embedded center coordinates as input, which are more global messages, and together with the branch takes in more local and detailed messages from neighbor points. Integrated these two different branches together, the structure is capable of mixing information of different scales. Thus the final output of the encoder then coped with a max or mean pooling layer, is actually a multi-scale feature extractor of point cloud, holding the same effect of both utilize and . This multi-scale feature, rich in semantic information, enables the model to understand local features and at the same time to accept global location information.
| Method | Reference | Params. (M) | ScanObjectNN | ModelNet40 | |||
| OBJ_BG | OBJ_ONLY | PB_T50_RS | 1k P | 8k P | |||
| Supervised Learning Only | |||||||
| PointNet (Qi et al. 2017a) | CVPR 17 | 3.5 | 73.3 | 79.2 | 68.0 | 89.2 | 90.8 |
| PointNet++ (Qi et al. 2017b) | NeurIPS 17 | 1.5 | 82.3 | 84.3 | 77.9 | 90.7 | 91.9 |
| DGCNN (Wang et al. 2019) | TOG 19 | 1.8 | 82.8 | 86.2 | 78.1 | 92.9 | - |
| PCT (Guo et al. 2021) | ICCV 21 | 2.88 | - | - | - | 93.2 | - |
| PointMLP (Ma et al. 2022) | ICLR 22 | 12.6 | - | - | 85.40.3 | 94.5 | - |
| PointNeXt (Qian et al. 2022) | NeurIPS 22 | 1.4 | - | - | 87.70.4 | 94.0 | - |
| with Self-Supervised Representation Learning (Full) | |||||||
| OcCo (Wang et al. 2021) | ICCV 21 | 22.1 | 84.85 | 85.54 | 78.79 | 1k | - / 92.1 |
| Point-BERT (Yu et al. 2022) | CVPR 22 | 22.1 | 87.43 | 88.12 | 83.07 | 93.2 | 93.8 |
| Point-MAE (Pang et al. 2022) | ECCV 22 | 22.1 | 90.02 | 88.29 | 85.18 | 93.8 | 94.0 |
| Point-M2AE (Zhang et al. 2022) | NeurIPS 22 | 15.3 | 91.22 | 88.81 | 86.43 | 94.0 | - |
| ACT (Dong et al. 2023) | ICLR 23 | 22.1 | 93.29 | 91.91 | 88.21 | 93.7 | 94.0 |
| VPP (Qi et al. 2023b) | NeurIPS 23 | 22.1 | 93.11 | 91.91 | 89.28 | 94.1 | 94.3 |
| I2P-MAE (Zhang et al. 2023c) | CVPR 23 | 15.3 | 94.15 | 91.57 | 90.11 | 94.1 | - |
| ReCon (Qi et al. 2023a) | ICML 23 | 44.3 | 95.18 | 93.29 | 90.63 | 94.1 | 94.3 |
| with Parameter-Efficient Supervised Finetuning | |||||||
| Point-MAE (Pang et al. 2022) (Baseline) | ECCV 22 | 22.1 | 90.02 | 88.29 | 85.18 | 93.8 | 94.0 |
| + IDPT∗ (Zha et al. 2023) | ICCV 23 | 1.7 | 92.94/92.38 | 92.60/91.33 | 88.34/88.09 | 93.64/93.02 | 93.88/93.67 |
| + DAPT∗ (Zhou et al. 2024) | CVPR 24 | 1.1 | 92.43/91.61 | 91.91/91.31 | 88.27/87.68 | 92.99/92.83 | 93.27/93.05 |
| + PPT (Ours) | - | 1.1 | 93.63/92.99 | 92.60/92.43 | 89.00/88.41 | 93.68/93.18 | 93.88/93.51 |
| ReCon (Qi et al. 2023a) (Baseline) | ICML 23 | 22.1 | 94.32 | 92.77 | 90.01 | 92.5 | 93.0 |
| + IDPT∗ (Zha et al. 2023) | ICCV 23 | 1.7 | 93.46/92.88 | 91.74/91.37 | 88.13/87.76 | 93.64/93.28 | 93.64/93.52 |
| + DAPT∗ (Zhou et al. 2024) | CVPR 24 | 1.1 | 93.63/93.12 | 92.43/91.63 | 89.31/88.77 | 93.27/92.94 | 93.07/92.79 |
| + PPT (Ours) | - | 1.1 | 95.01/94.09 | 93.28/93.23 | 89.52/88.97 | 93.76/93.41 | 93.84/93.66 |
Better way of prompt and adapter tuning?
With point cloud as the only modality of input, it is difficult for 3D parameter-efficient fine-tuning tasks to get prompts that are rich in linguistic significance, like a well-designed prefix prompt in natural language, which can remarkably boost the performance of generative tasks in text modality (Lester, Al-Rfou, and Constant 2021). VPT (Jia et al. 2022b) introduce a set of trainable embeddings into the input of the Transformer to solve this problem in the vision domain, and both DAPT (Zhou et al. 2024) and IDPT (Zha et al. 2023) adopt a similar manner. Simultaneously considering Adapter and Prompt, they both generate extra prompt tokens from an adapter-structure network and add these prompt tokens into Transformer input. By doing so they acquire dynamic prompts rather than static ones in VPT. As the static prompts always need a process of manual selection, the generated dynamic prompts are a compromise of this process. Nevertheless, these kinds of prompts are actually an abstraction of the existing feature embedding. Thus the input of Transformer layers with extra prompt inserted and the Transformer output can be formulated as:
Where means class token, represents the extra prompts abstracted from patch embedding . attn and F represent the FFN and attention layer inside the Transformer blocks. The and represent trainable and frozen parameters, respectively. Thus the information of prompts is highly homogeneous, which is an aggregation of existing tokens, rather than the newly added language prompts in the NLP domain. We then think about whether there are some other forms of prompts that can both be dynamically adjusted and in the meantime be of physical significance And it comes to our mind that the patch tokens themselves, which are down-sampled and clustered, can be used as extra tokens for prompting. Different from prompts that are yielded from patch token embedding, we unfreeze a part of the patch encoder to introduce variety into patch token embedding. In that case, with this form of prompts, the input of Transformer layers can be formulated as:
Here and mean different adapter layers for dynamic adjustment. Thus the output of the Transformer is:
| Method | Reference | 5-way | 10-way | ||
| 10-shot | 20-shot | 10-shot | 20-shot | ||
| DGCNN (Wang et al. 2019) | TOG 19 | 31.6 2.8 | 40.8 4.6 | 19.9 2.1 | 16.9 1.5 |
| OcCo (Wang et al. 2021) | ICCV 21 | 90.6 2.8 | 92.5 1.9 | 82.9 1.3 | 86.5 2.2 |
| with Self-Supervised Representation Learning (Full) | |||||
| OcCo (Wang et al. 2021) | ICCV 21 | 94.03.6 | 95.92.3 | 89.45.1 | 92.44.6 |
| Point-BERT (Yu et al. 2022) | CVPR 22 | 94.6 3.1 | 96.3 2.7 | 91.0 5.4 | 92.7 5.1 |
| Point-MAE (Pang et al. 2022) | ECCV 22 | 96.3 2.5 | 97.8 1.8 | 92.6 4.1 | 95.0 3.0 |
| Point-M2AE (Zhang et al. 2022) | NeurIPS 22 | 96.8 1.8 | 98.3 1.4 | 92.3 4.5 | 95.0 3.0 |
| ACT (Dong et al. 2023) | ICLR 23 | 96.8 2.3 | 98.0 1.4 | 93.3 4.0 | 95.6 2.8 |
| VPP (Qi et al. 2023b) | NeurIPS 23 | 96.9 1.9 | 98.3 1.5 | 93.0 4.0 | 95.4 3.1 |
| I2P-MAE (Zhang et al. 2023c) | CVPR 23 | 97.0 1.8 | 98.3 1.3 | 92.6 5.0 | 95.5 3.0 |
| ReCon (Qi et al. 2023a)(baseline) | ICML 23 | 97.3 1.9 | 98.9 1.2 | 93.3 3.9 | 95.8 3.0 |
| with Parameter-Efficient Supervised Finetuning | |||||
| ReCon w/ IDPT∗ (Zha et al. 2023) | ICCV 23 | 96.92.4 | 98.30.7 | 92.84.0 | 95.53.2 |
| ReCon w/ DAPT∗ (Zhou et al. 2024) | CVPR 24 | 95.62.8 | 97.71.6 | 91.94.1 | 94.63.5 |
| ReCon w/ PPT (Ours) | - | 97.0 2.7 | 98.71.6 | 92.25.0 | 95.62.9 |
represents the added trainable adapter for more thorough feature aggregation and adjustment. In this way, both the extra sampled patch token prompts and the original patch token embedding are altered for information integration. On this manner of prompts and adapters, combined with the aforementioned positional encoding, we carried out extensive experiments, and our method achieved excellent results in all of these experiments, which will be discussed in subsequent chapters.
Pipeline of PPT
The pipeline structure of our PPT is shown in Figure 3(a). The patch encoder of the pre-trained model is already empowered to abstract a high concentration of semantic information from the input point cloud, so we unfreeze part of the patch encoder to provide varieties for patch tokens and prompt tokens. Briefly, we resample the input point clouds as prompts and concatenate them with the input patch tokens before sending them into the Transformer, and trainable positional embedding layers are concatenated together as the positional information for both the prompts and the patch tokens. Before inputting into each Transformer layer, a patch adapter layer and a prompt adapter layer are adopted as a dynamic adjustment module. There are also adapter layers inside every Transformer block, specifically between the attention layer and the FFN. As the pre-trained parts of the Transformer layers are untrainable during the tuning process, an adapter inside the block effectively integrates information and dynamically adjusts it to improve performance.
Experiments
3D Real-World Object Recognition
ScanObjectNN (Uy et al. 2019) is one of the most challenging 3D datasets, which covers 15K real-world objects from 15 categories. For a fair comparison, we report the results with and without the voting strategy (Liu et al. 2019) separately. The results of object classification tasks of ScanObectNN (OBJ_BG, OBJ_ONLY, PB_T50_RS) are shown in Table 1. For each variant, we conducted experiments 8 times with different random seeds and reported the best and the average scores, and it is observed that: (i)Holistically, with only 1.05M, 4.7% parameters of Full fine-tuning paradigm, the performance of our block is improved by +0.7% and +11.3% compared with that of standard Transformer under Full tuning protocol on ReCon and Point-MAE, respectively. (ii) Compared with the current best methods IDPT and DAPT, our PPT also gets better results than both of them. Specifically, we achieve 2.4% and 4.5% performance improvement compared with DAPT and IDPT respectively. That is, with PPT on both ReCon and Point-MAE achieve state-of-the-art performance.
3D Synthetic Object Recognition
ModelNet (Wu et al. 2015) is one of the most classical datasets for synthetic 3D object recognition. It contains 12K meshed 3D CAD objects of 40 (ModelNet40) or 10 (ModelNet10) categories. We conducted the evaluation on the ModelNet40 dataset, including fine-tuning and few-shot learning. During training and testing, we use Scale&Translate as data augmentation in training following (Qi et al. 2017a, b). The results are shown in Table 1 and Table 2, respectively. We report two different settings with 1k points and 8k points respectively. It can observed that (i)With only a few trainable parameters, compared with Full tuning protocol, ReCon gains 2.1% performance improvement with our PPT. Point-MAE also gets results on par with the Full tuning with less than 5% of the parameters trainable. (ii)Compared with the current best methods IDPT and DAPT, with our PPT, we achieve 1.26% and 0.32% improvements higher than these two respectively. Point-MAE with PPT achieves the best result on ModelNet40 with 93.88% accuracy. And ReCon with PPT gains +1.3% and +0.8% performance improvements, showing the effectiveness of our positional prompt tuning method.
| Methods | Reference | #TP (M) | Cls. mIoU (%) | Inst. mIoU (%) |
| Supervised Learning Only | ||||
| PointNet (Qi et al. 2017a) | CVPR 17 | - | 80.39 | 83.7 |
| PointNet++ (Qi et al. 2017b) | NeurIPS 17 | - | 81.85 | 85.1 |
| DGCNN (Wang et al. 2019) | TOG 19 | - | 82.33 | 85.2 |
| Self-Supervised Representation Learning (Full fine-tuning) | ||||
| OcCo (Wang et al. 2021) | ICCV 21 | 27.09 | 83.42 | 85.1 |
| Point-BERT (Yu et al. 2022) | CVPR 22 | 27.09 | 84.11 | 85.6 |
| Point-MAE (Pang et al. 2022) | ECCV 22 | 27.06 | 84.19 | 86.1 |
| ACT(Dong et al. 2023) | ICLR 23 | 27.06 | 84.66 | 86.1 |
| ReCon(Qi et al. 2023a) | ICML 23 | 27.06 | 84.66 | 86.4 |
| with Parameter-Efficient Supervised Finetuning | ||||
| Point-MAE (Pang et al. 2022) (baseline) | ECCV 22 | 27.06 | 84.19 | 86.1 |
| + IDPT (Zha et al. 2023) | ICCV 23 | 5.69 | 83.79 | 85.7 |
| + DAPT (Zhou et al. 2024) | CVPR 24 | 5.65 | 84.01 | 85.7 |
| + PPT (ours) | - | 5.62 | 84.07 | 85.7 |
| ReCon (Qi et al. 2023a) (baseline) | ICML 23 | 27.06 | 84.52 | 86.1 |
| + IDPT (Zha et al. 2023) | ICCV 23 | 5.69 | 83.66 | 85.7 |
| + DAPT (Zhou et al. 2024) | CVPR 24 | 5.65 | 83.87 | 85.7 |
| + PPT (ours) | - | 5.62 | 84.23 | 85.6 |
Few-shot Learning
We conducted experiments of few-shot learning on the ModelNet40 (Wu et al. 2015) dataset. We considered settings of n-way k-shot manner following previous works (Pang et al. 2022; Qi et al. 2023a), where n 5, 10 and k 10, 20. As shown in Table 2, it can be observed that our PPT can achieve performance enhancement in most cases compared with IDPT and DAPT, which also indicates the effectiveness of our method.
3D Part Segmentation
To evaluate the geometric understanding performance within objects, we conducted the part segmentation experiment on ShapeNetPart (Yi et al. 2016). Specifically, we concatenate the cross-modal feature into the global feature and use the same segmentation head as Point-MAE for a fair comparison. From Table 3, it can be observed that our method gets results on par with from scratch baseline on both Cls. mIoU and Inst. mIoU. Besides, PPT also outperforms the PEFT counterparts IDPT (Zha et al. 2023) and DAPT (Zhou et al. 2024) 0.06% and 0.46% Cls. mIoU, respectively. It shows that the cross-modal global knowledge from cross-modal pretraining can still play a certain role in part segmentation.
| Methods | Params. (M) | OBJ_BG | OBJ_ONLY | PB_T50_RS |
| PointNet | 3.5 | 73.3 | 79.2 | 68.0 |
| PointNet++ | 1.5 | 82.3 | 84.3 | 77.9 |
| Point-BERT | 22.1 | 87.43 | 88.12 | 83.07 |
| PosNet (Ours) | 0.8 | 90.02 | 89.28 | 84.41 |
Discussions
Position is all your need in 3D Transformer?
Positional encoding is particularly crucial in 3D Transformers. We have demonstrated that combining MLP-based positional encoding with local patch tokens effectively functions as both global and local point cloud feature extractors (e.g., PointNet). A bold hypothesis is whether, without the Transformer blocks, positional encoding tokens alone can achieve comparable or even superior performance.
Figure 4 illustrates this very simple architecture, and its classification performance on ScanObjectNN is shown in Table 4. With only 0.8M parameters, PosNet achieved remarkable performance, further validating our theoretical assumption that the global understanding provided by positional encoding is crucial in 3D Transformers.
Conclusions
In this study, we rethink the role of positional encoding in 3D representation learning and fine-tuning. We argue that using positional encoding in point Transformer-based methods serves to aggregate multi-scale features of point clouds. Additionally, we explore parameter-efficient fine-tuning (PEFT) through the lens of prompts and adapters, introducing a straightforward yet effective method called PPT for point cloud analysis. PPT incorporates increased patch tokens and trainable positional encoding while keeping most pre-trained model parameters frozen. Extensive experiments validate that PPT is both effective and efficient.
Appendix A Additional Experiments
Implement Details
To ensure a fair comparison, we employed identical experimental settings to the default fine-tuning method for each baseline. This entails freezing the weights of the pre-trained point cloud backbone and solely updating the newly inserted parameters during training. All experiments are conducted on a single GeForce RTX 3090.
Ablation Studys
We have done several ablation experiments on our method. In Table 7 we replace the positional embedding layer in the PPT with several classic forms. The means position encoding is frozen during the tuning process. As we can see in the table, the results show that such MLP structures as position embedding layers, as we mentioned before, perform better than others, which also proves that the point Transformer with MLP embedding layers serves as a multi-scale feature extractor. And it’s surprising that tuning with only the first layer receives MLP-embedded positional information can achieve relatively high performance, even better than all other forms. We also set an ablation study on the number of patch token groups K. Specifically, we set the number K range from 1 to 3, and K=2 turns out to be the best choice, demonstrating the efficiency of the extra prompt tokens, and also, too many prompt token groups will interference the process of fine-tuning. Nevertheless, the result of K=3 which is relatively higher than the result of K=1 also shows the effectiveness of the extra token prompts edgewise.
| Classification | Few-Shot | Segmentation | ||
| Config | ScanObjectNN | ModelNet | ModelNet | ShapeNetPart |
| optimizer | AdamW | AdamW | AdamW | AdamW |
| learning rate | 2e-5 | 1e-5 | 5e-4 | 2e-4 |
| weight decay | 5e-2 | 5e-2 | 5e-2 | 5e-2 |
| learning rate scheduler | cosine | cosine | cosine | cosine |
| training epochs | 300 | 300 | 150 | 300 |
| warmup epochs | 10 | 10 | 10 | 10 |
| batch size | 32 | 32 | 32 | 16 |
| drop path rate | 0.3 | 0.3 | 0.2 | 0.1 |
| number of points | 2048 | 1024/8192 | 1024 | 2048 |
| number of point patches | 128 | 64/512 | 64 | 128 |
| point patch size | 32 | 32 | 32 | 32 |
| augmentation | Rotation | Scale&Trans | Scale&Trans | - |
| GPU device | RTX 3090 | RTX 3090 | RTX 3090 | RTX 3090 |
| K | #TP (M) | PB_T50_RS |
| 1 | 0.9M | 87.40 |
| 2 | 1.1M | 89.52 |
| 3 | 1.4M | 88.06 |
Appendix B Visualization
t-SNE Visualization
In Figure 5, we report the t-SNE visualization results of IDPT(Zha et al. 2023), DAPT(Zhou et al. 2024), and our PPT. Here we report the point cloud feature manifold visualization results on the ScanObjectNN PB_T50_RS dataset. It can be observed that our method, with fewer or on par number of parameters, gets more compact clusters and more separate cluster centers, which means our method gets better results than the parameter-efficient full fine-tuning counterparts.
Attention Visualization
We also report the attention score visualization of either position embedding or patch tokens as the input respectively. Taking the 3D coordinates of patch centers as the input of position embedding, the attention scores of the position embedding input are more holistic, as shown in the top row of Figure 6. That means the position embedding features are concerned with more global information. Meanwhile, the patch token embedding, with all neighbor points around the cluster centers as input, their attention gathered together at some points, as the local counterpart of position embedding inputs. Both concerning global and local features, the point Transformer architectures are deal with complex tasks with a simple MLP as the position encoding module and even get better results than adding those well-designed position embedding methods together, as reported in Table 7.
Appendix C Limitations & Future Work
Although we want to underscore the significance of positional encoding within 3D Transformers and have experimentally validated the use of positional encoding as a form of prompt tuning in classic 3D self-supervised pre-training works such as Point-MAE (Pang et al. 2022) and ReCon (Qi et al. 2023a), we have yet to conduct experimental validation on larger datasets, such as Objaverse (Deitke et al. 2023). Our primary future work will be to experiment with positional prompt tuning on a wider range of 3D foundational models, such as OpenShape (Liu et al. 2023) and ReCon++ (Qi et al. 2024), and to expand its application to a broader spectrum of tasks, including zero-shot learning (Xue et al. 2023; Liu et al. 2023; Qi et al. 2023a, 2024), 3D AIGC (Nichol et al. 2022; Qi et al. 2023b), and 3D LLMs (Qi et al. 2024; Dong et al. 2024), among others (Dong et al. 2022; Peng et al. 2024; Zhang et al. 2023b).